# Radial Profiles of Radio Halos in Massive Galaxy Clusters: Diffuse Giants Over 2 Mpc

### v1

@@ 3 Radio continuum maps of individual clusters
Clearly, in all cases, the radio emission is brightest at the halo core and decreases with radial distance. Radio halos with LLS$>$2 Mpc have been mainly detected at low frequencies, especially with observations $\leq 150$ MHz (e.g.; Botteon et al., 2018; Shweta et al., 2020; Botteon et al., 2022b; Bonafede et al., 2022; Rajpurohit et al., 2021b; Bruno et al., 2023). Our results highlight the emergence of radio halos with large spatial extents even at high frequencies ($>$700 MHz) and demonstrate that their observed size is determined by the depth of the radio observations and uv-coverage, and their radio brightness. The detection of radio halos with large extents ($>$ 2Mpc) shows the presence of relativistic electrons and magnetic field beyond $R_{500}$. It is, therefore, important to establish their observable quantities that allow a discrimination between these observed sources with others (e.g., mega halos).
@@ 4 Radial properties of radio halos
A handful of radio halos with remarkably large extents ($\geq 2$ Mpc) are currently known, see Table 4. These systems are unique targets to test whether the physics of non thermal components at such large scales is different from that at cluster-core scales and to probe to origin of the radio emission. The detection of halo emission, on such large scales, at multiple frequencies, in the five clusters analyzed in this work provides a unique opportunity to investigate their properties and to shed light on the underlying particle acceleration mechanisms.
@@ 4.1 Radial surface brightness profiles
where $I(r)$ is the surface brightness at radius $r$, $I_{0}$ is the central radio brightness, and $r_{e}$ is the e-folding radius. We emphasize that this model is not physically motivated. Although it has historically provided a reasonable description of halo profiles (e.g., Orrú et al., 2007; Murgia et al., 2009; Vacca et al., 2011; Cuciti et al., 2021), the detection of substructures and multiple components in halos with new generation telescopes has highlighted its limitations (Botteon et al., 2022a; Rajpurohit et al., 2021b; Botteon et al., 2023; Sikhosana et al., 2023). Moreover, this model grossly oversimplifies the cluster morphology; merging systems can exhibit highly extended and irregular structures, see e.g., the PLCK G287.0+32.9 and Bullet halos. Nevertheless, this model is useful for comparative studies of different kinds of halos and, therefore, we adopt it in this paper to facilitate the comparison.
For each cluster, we first computed the average radio brightness within concentric annuli centered on the peak of the radio halo emission, with the width of each annulus set to half the full width at half maximum of the beam size, see Table 5. The uncertainties in the profiles are estimated by accounting for systematic uncertainties due to flux scale/discrete sources subtraction (assumed to be 5-10%) and statistical uncertainties associated with the rms level of an image. We included only those annuli where the average surface brightness is $\geq 3\sigma_{\rm rms}$. The discrete, unrelated sources (namely, compact radio sources, radio galaxies, relics) were masked out from each map, as discussed in detail in Section 5.
@@
In Figures 7 and 8, we show the resulting radial surface brightness profiles. For the first time, we present radial surface brightness profiles of radio halos out to $R_{500}$ as a function of frequency. Each point in these plots represents the mean radio brightness measured within each concentric circular annulus. The best fit parameters for each cluster are summarized in Table 5. Remarkably, in all cases, the halo emission can be described by a single exponential fit, despite the simplistic assumption of a symmetrical morphology and the highly disturbed dynamical nature of the clusters, which contain several substructures.
The radial profiles of the PLCK G287.0+32.9 halo are well described by a single-component exponential fit at 350 MHz, 815 MHz, 1.28 GHz, and 2.4 GHz (see Figure 7 top-left panel). The best-fit e-folding radius varies with observing frequency, being smaller at lower frequencies. This is simply because the fainter outer regions are not detected at high-signal to ratio at 350 MHz due to the lower sensitivity and missing short baselines of the uGMRT Band3 observations compared to the MeerKAT. However, between MeerKAT 815 MHz and 1.28 GHz the total extent of the halo is similar and the corresponding best fit e-folding radii indicate a clear radial steepening of the spectral index.
@@
The halo in the Bullet Cluster also shows a kink in its radial radio profile (Figure 7, bottom-left panel), which coincides with the location of the bow shock detected via X-ray observations (Markevitch et al., 2002). This region is not masked as no relic-like emission is observed in the radio maps. Therefore, the kink is associated with the shock, similar to what is observed in Abell 2744, possibly suggesting a connection between shocks and turbulence. The radio profiles of both the Abell 2744 and Bullet Cluster halos show larger e-folding radii at higher frequencies (when comparing the same extent of the halo emission), hinting at flatter spectral indices in the outermost regions. This trend is also evident in the radial profiles, with the 815 MHz profile steepening more rapidly at large radii than the 1.28 GHz.
@@
In summary, despite the complex morphologies of the radio emission, the radial surface brightness profiles of the five halos with extents $>$2 Mpc, analyzed here, are consistent with other known halos ($<$2 Mpc). Thanks to the sensitivity of new generation telescopes, these five halos are detected at very large distances. Our sample consists of five radio halos but we do not find evidence of a shallower second component in the outer regions as reported by Cuciti et al. (2022) for four mega-halos.
@@ 4.2 Radial spectral index profile
The radio observations presented in this study were obtained using different interferometers, each characterized by distinct uv-coverages. Consequently, careful consideration is required when comparing flux density measurements/extracting spectral index values for extended emission. To obtain the spectral index, we created maps using Briggs weighting with a robust parameter of $-0.5$. To ensure a consistent flux density distribution across all observed frequencies, we applied a common inner uv-cut at $200\lambda$, corresponding to the well-sampled shortest baseline of the uGMRT data. For MACS J0717+3745, we use the published 20″ and 75″ images, respectively.
@@
The resulting azimuthally averaged spectral index profiles, centered on the point of peak surface brightness, are shown in Figure 9. The halos in PLCK G287.0+32.9 and MACS J0717.5+3745 (previously reported by Rajpurohit et al. 2021b) show clear evidence of radial spectral steepening, also reported with other halos, e.g., Coma (Bonafede et al., 2022). The low frequency radial spectral index in the PLCK G287.0+32.9 halo steepens from $-0.9$ in the inner most region ($<$250 kpc) to $-1.8$ at a distance of 1.3 Mpc from the cluster center. The same trends are observed toward high frequencies with steepening up to $-2.5$ in the outermost areas.
For the halo in Abell 2744, the low frequency radial spectral index profile remains approximately constant (excluding the innermost region, $<$150 kpc) at around $-1.15$ out to 800 kpc from the cluster center, consistent with the findings of Orrú et al. (2007). However, the high-frequency ($1.5-3$ GHz) spectral index profile shows a clear spectral steepening in the outer regions, also reported by Pearce et al. (2017). Similarly the MACS J0717.5+3745 halo also shows a radial steepening from $-1$ in the innermost region to $-1.8$ at 900 kpc from the cluster center. We emphasize that in all cases, the outermost regions of the halo is steeper as also reported for mega halos.
For the halos in PLCK G287.0+32.9 and Abell 2744, the spectral index distribution is flatter at low frequencies and steepens toward higher frequencies, indicating clear spectral curvature. This is not surprising, as reported by Rajpurohit et al. (2021a, 2023) the radio halos exhibit different spectral indices and curvature distributions. The radial spectral index steepening observed in the these halos is consistent with the general expectations of the turbulent reacceleration models (Brunetti & Jones, 2014).
@@
Cluster $S_{\rm measured}$ $S_{\rm fitted}$ Measured radio power (lower limit) Fitted Radio power
1.5 GHz (mJy) mJy $\rm 10^{25}W\,Hz^{-1}$ $\rm 10^{25}W\,Hz^{-1}$
@@
Abell 2142 $26\pm 2$ $44\pm 5$ $0.05\pm 0.01$ $0.08\pm 0.01$
Notes: Column 1: name of the cluster hosting the halo, Col. 2: Measured flux density of the halo at 1.5 GHz, Col. 3: Fitted flux density of the halo, Col. 4: Radio power obtained from flux density reported in Col 2, Col. Radio power obtained from flux density reported in Col 3
@@ 4.3 Radio emissivity and radio power
Radio halos are observed to exhibit a relatively narrow distribution of radio emissivity, with a characteristic value of approximately $\rm 10^{-42}\,ergs^{-1}\,cm^{-3}\,Hz^{-1}$ (e.g., Murgia et al., 2009). This is based on the assumption of a homogeneous/filling factor=1 emitting volume. The emissivity of mega-halos (second component) is reported to be approximately $10^{-44}\,\rm ergs^{-1}\,cm^{-3}\,Hz^{-1}$, about a factor of 20-25 lower than that of classical radio halos (Cuciti et al., 2022).
@@
where $I_{0}$ is the central surface brightness in $\mu$Jy/arcsec2, $r_{e}$ is the e-folding radius in kpc, and $\alpha$ is the integrated radio spectral index. The term $(1+z)^{3-\alpha}$ takes into account of both the $k$-correction and cosmological dimming of the surface brightness with redshift. The values of $I_{0}$ and $r_{e}$ are obtained from fitting an exponential model to the radial surface brightness profiles. The radio emissivities for the halos in our sample are reported in Table 5.
We find that the observed halos display a comparable emissivity of approximately $10^{-42}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$. In Figure 10, we present the $I_{0}-r_{e}$ plane at 1.4 GHz for our sample (red points), along with a comparison to a sample of 14 known halos from (Murgia et al., 2009; Vacca et al., 2011; Murgia et al., 2024). We note that Murgia et al. (2024) also masked unrelated sources, making the results directly comparable. Despite their different LLS, we find that the emissivities of halos with LLS>2 Mpc are remarkably similar to those of classical halos (Murgia et al., 2024). Moreover, the e-folding radius of these five halos is only about 50 kpc larger than the average e-folding radius reported by Murgia et al. (2024)
The halo in Abell 2142 is different as its emissivity is about a factor of 10 lower than the other halos with LLS$>2$ Mpc, namely $\langle J\rangle=(1.43\pm 0.09)\times 10^{-43}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$ (estimated using an integrated spectral index of $-1.2$, i.e., mean spectral index of the three components). The prototype radio halo in Coma has the lowest emissivity observed so far at 1.4 GHz, with $\langle J\rangle\simeq 5\times 10^{-44}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$ (Murgia et al., 2024). We note that the Coma and Abell 2142 halos are peculiar, although both are nearby clusters and their radii are comparable to other halos with large extents, their central brightnesses are about an order of magnitude fainter (see Figure 10). The lower emissivities of the Abell 2142 and Coma halos suggests that halos with such low emissivity may not be uncommon.
@@
where $D_{L}$ is the luminosity distance to the cluster, $S_{\nu}$ is the flux density at a frequency $\nu$. We measured the flux density of these halos from 1.4 GHz maps, see Table 6. The derived radio powers, at 1.4 GHz, are $P_{\rm PLCK\,G287}=(1.3\pm 0.1)\times 10^{25}\,\mathrm{W\,Hz^{-1}}$, $P_{A2744}=(1.4\pm 0.1)\times 10^{25}\,,\mathrm{W\,Hz^{-1}}$ (Rajpurohit et al., 2021a; Pearce et al., 2017), $P_{\rm MACS\,J0717}=(2.2\pm 0.3)\times 10^{25}\,\mathrm{W\,Hz^{-1}}$ (Rajpurohit et al., 2021b), $P_{\rm Bullet}=(2.7\pm 0.4)\times 10^{25}\,\mathrm{W\,Hz^{-1}}$, $P_{A2142}=(5.2\pm 1.3)\times 10^{23}\,\mathrm{W\,Hz^{-1}}$ (Riseley et al., 2024). These values represent lower limits, as part of the halo emission is excluded/masked due to the presence of discrete sources.
Figure 11: Radial profiles of the halos in Abell 2142 (left) and PLCK G287.0+32.9 (right). Profiles fitted with a single-component exponential model are shown with dashed lines. The chosen sectors are shown in Figure 6. Left: Radial profile of the Abell 2142 halo at 144 MHz, extracted from an image where discrete sources were masked based on the LOFAR HBA map. The blue data points/line show the residual radial profile obtained subtracting the $uv$-subtracted discrete sources profile from the masked out sources profile. Top right: PLCK G287.0+32.9 halo profiles at 815 MHz and 350 MHz obtained after subtracting discrete sources from the $uv$-data. Both plots show a second outer component due to incomplete discrete sources subtraction.
@@
[…]
where $r_{e}$ and $I_{0}$ are in units of arcseconds and $\rm mJy\,arcsec^{-2}$, respectively, and $f\approx 0.8$ accounts for the enclosed flux within $3r_{e}$. The fitted flux densities and the corresponding radio powers are summarized in Table 6. The fitted and measured flux densities of the radio halos in Abell 2744, MACS J0717+3745, and the Bullet Cluster are consistent. However, the fitted flux densities for the halos in PLCK G287.0+32.9 and Abell 2142 are lower than the measured values by a factor of two. Murgia et al. (2009) reported that a tight relation between the total radio power of halos scales and e-folding radius, i.e., $P_{1.4}\propto r_{e}^{3}$. The halos analyzed in this work are consistent with this relation and align well with previously known classical halos. Moreover, powerful radio halos are expected to be more extended, corresponding to larger e-folding radii (Cassano et al., 2007; Murgia et al., 2009). This implies that powerful halos should be detectable out to larger radii.
The lower limits and the fitted radio power imply that, excluding the halo in Abell 2142, the remaining systems lie above the established scaling relation between cluster mass and radio power observed for halos at 1.4 GHz. In contrast, the halo in Abell 2142 lies below this relation, as reported by Riseley et al. (2024); Bruno et al. (2023). This also implies that radio powers of halos reported in the literature may be either underestimated (Shimwell et al., 2014; Bonafede et al., 2014) or overestimated (van Weeren et al., 2009; Bonafede et al., 2009; Venturi et al., 2017) due to poor sensitivity and resolution. This in turn impacts the established $P\text{?}M_{500}$ scaling relations of radio halos.

### v2

@@ 3 Radio continuum maps of individual clusters
Clearly, in all cases, the radio emission is brightest at the halo core and decreases with radial distance. Radio halos with LLS>>2 Mpc have been mainly detected at low frequencies, especially with observations $\leq 150$ MHz (e.g., Botteon et al., 2018; Shweta et al., 2020; Botteon et al., 2022b; Bonafede et al., 2022; Rajpurohit et al., 2021b; Bruno et al., 2023). Our results highlight the emergence of radio halos with large spatial extents even at high frequencies (>>700 MHz) and demonstrate that their observed size is determined by the depth of the radio observations and uv-coverage, and their radio brightness. The detection of radio halos with large extents (>> 2Mpc) indicates the presence of relativistic electrons and magnetic fields beyond $R_{500}$. It is, therefore, important to establish the observable quantities that allow a discrimination between these sources and others (e.g., mega halos).
@@ 4 Radial properties of radio halos
A handful of radio halos with remarkably large extents ($\geq 2$ Mpc) are currently known, see Table 5. These systems are unique targets to test whether the physics of non-thermal components at such large scales is different from that at cluster-core scales and to probe the origin of the radio emission. The detection of halo emission, on such large scales, at multiple frequencies, in the five clusters analyzed in this work, provides a unique opportunity to investigate their properties and to shed light on the underlying particle acceleration mechanisms.
@@ 4.1 Radial surface brightness profiles
where $I(r)$ is the surface brightness at radius $r$, $I_{0}$ is the central radio brightness, and $r_{e}$ is the e-folding radius. We emphasize that this model is not physically motivated. Although it has historically provided a reasonable description of halo profiles (e.g., Orrú et al., 2007; Murgia et al., 2009; Vacca et al., 2011; Cuciti et al., 2021), the detection of substructures and multiple components in halos with new generation telescopes has highlighted its limitations (Botteon et al., 2022a; Rajpurohit et al., 2021b; Botteon et al., 2023; Sikhosana et al., 2023). Moreover, this model grossly oversimplifies the cluster morphology; merging systems can exhibit highly extended and irregular structures, for example, the PLCK G287.0+32.9 and Bullet halos. Nevertheless, this model is useful for comparative studies of different kinds of halos and, therefore, we adopt it in this paper to facilitate the comparison.
For each cluster, we first computed the average radio brightness within concentric annuli centered on the peak of the halo emission, with the width of each annulus set to half the full width at half maximum of the beam size, see Table 6. The circular annuli used are shown in Figure 7. The uncertainties in the profiles are estimated by accounting for systematic uncertainties due to flux scale/discrete sources subtraction (assumed to be 5-10%) and statistical uncertainties associated with the rms level of an image. We assumed a flux density scale error of 10% for LOFAR, uGMRT Band3 (Shimwell et al., 2022; Chandra & Kanekar, 2017), 8% for uGMRT Band4 (Chandra & Kanekar, 2017), and 5% for VLA and MeerKAT (Perley & Butler, 2013). We note that the flux density scale was verified by checking the spectra of compact sources in each field, we also refer to van Weeren et al. (2017); Pearce et al. (2017); Rajpurohit et al. (2021a, b, c) for details. We included only those annuli where the average surface brightness is $\geq 3\sigma_{\rm rms}$. The discrete, unrelated sources (namely, compact radio sources, radio galaxies, relics) were masked out from each map, as discussed in detail in Section 5.
@@
In Figures 8 and 9, we show the resulting radial surface brightness profiles. For the first time, we present radial surface brightness profiles of radio halos out to $R_{500}$ as a function of frequency. Each point in these plots represents the mean radio brightness measured within each concentric circular annulus. The best fit parameters for each cluster are summarized in Table 6. Remarkably, in all cases, the halo emission can be described by a single exponential fit, despite the simplistic assumption of a symmetrical morphology and the highly disturbed dynamical nature of the clusters, which contain several substructures.
The radial profiles of the PLCK G287.0+32.9 halo are well described by a single-component exponential fit at 350 MHz, 815 MHz, 1.28 GHz, and 2.4 GHz (see Figure 8 top-left panel). The best-fit e-folding radius varies with observing frequency, being smaller at lower frequencies. This is simply because the fainter outer regions are not detected at high signal-to-noise ratio at 350 MHz due to the lower sensitivity and missing short baselines of the uGMRT Band3 observations compared to the MeerKAT. However, between MeerKAT 815 MHz and 1.28 GHz, the total extent of the halo is similar and the corresponding best-fit e-folding radii indicate a clear radial steepening of the spectral index.
@@
The halo in the Bullet Cluster also shows a kink in its radial radio profile (Figure 8, bottom-left panel), which coincides with the location of the bow shock detected via X-ray observations (Markevitch et al., 2002). This region is not masked as no relic-like emission is observed in the radio maps. Therefore, the kink is associated with the shock, similar to what is observed in Abell 2744, possibly suggesting a connection between shocks and turbulence. The radial profiles of both the Abell 2744 and Bullet Cluster halos show larger e-folding radii at higher frequencies (when comparing the same extent of the halo emission), hinting at relatively flatter spectral indices in the outermost regions. This is also evident from the 815 MHz profile, which steepens more rapidly at large radii than the 1.28 GHz profile.
The 400 MHz and 144 MHz radial profiles of MACS J0717.5+3745 is shown in Figure 8, bottom-right panel. Despite the complex morphology of the halo emission, the profiles are well described by a single exponential fit. The halo is more extended at lower frequencies, therefore, the e-folding radius is larger at 144 MHz, suggesting the spectral steepening in the outermost regions. We note that the halo in MACS J0717.5+3745 also exhibits some filamentary features embedded in it (van Weeren et al., 2017; Rajpurohit et al., 2021c). However, those filaments are not associated with the halo emission (Rajpurohit et al., 2022a).
@@
In summary, despite the complex morphologies of the radio emission, the radial surface brightness profiles of the five halos with extents >>2 Mpc, analyzed here, are consistent with other known halos (<<2 Mpc). Thanks to the sensitivity of new generation telescopes, these five halos are detected at very large distances. Our sample consists of five radio halos but we do not find evidence of a shallower second component in the outer regions as reported by Cuciti et al. (2022) for four mega-halos.
@@ 4.2 Radial spectral index profile
The radio observations presented in this study were obtained using different interferometers, each characterized by distinct uv-coverages. Consequently, careful consideration is required when comparing flux density measurements/extracting spectral index values for extended emission. To obtain the spectral index, we created maps using Briggs weighting with a robust parameter of 0.5-0.5 To ensure a consistent flux density distribution across all observed frequencies, we applied a common inner uv-cut at $200\lambda$, corresponding to the well-sampled shortest baseline of the uGMRT data, see see Table 4 for imaging properties (IM5-IM8, IM11-IM12, and IM16-IM19). For Abell 2142, we use the published 75″ images. We note that the same images are used to estimate the radio power of halos.
@@
The resulting azimuthally averaged spectral index profiles, centered on the point of peak surface brightness, are shown in Figure 10. The halos in PLCK G287.0+32.9 and MACS J0717.5+3745 (previously reported by Rajpurohit et al. 2021b) show clear evidence of radial spectral steepening, also reported for other halos, e.g., Coma (Bonafede et al., 2022). The low frequency radial spectral index in the PLCK G287.0+32.9 halo steepens from 0.9-0.9 in the innermost region (<250<250 kpc) to 1.8-1.8 at a distance of 1.3 Mpc from the cluster center. The same trends are observed toward high frequencies with steepening up to 2.5-2.5 in the outermost areas.
For the halo in Abell 2744, the low frequency radial spectral index profile remains approximately constant (excluding the innermost region, <<150 kpc) at around 1.15-1.15 to 800 kpc from the cluster center, consistent with the findings of Orrú et al. (2007). However, the high-frequency (1.531.5-3 GHz) spectral index profile shows a clear spectral steepening in the outer regions, also reported by Pearce et al. (2017). Similarly, the MACS J0717.5+3745 halo also shows a radial steepening from 1-1 in the innermost region to 1.8-1.8 at 900 kpc from the cluster center. We emphasize that in all cases, the outermost regions of the halo are steeper, as also reported for mega halos.
For the halos in PLCK G287.0+32.9 and Abell 2744, the spectral index distribution is flatter at low frequencies and steepens toward higher frequencies, indicating clear spectral curvature. This is not surprising, as reported by Rajpurohit et al. (2021a, 2023), the radio halos exhibit different spectral indices and curvature distributions. The radial spectral index steepening observed in these halos is consistent with the general expectations of the turbulent reacceleration models (Brunetti & Jones, 2014).
@@
Cluster $S_{\rm measured}$ $S_{\rm fitted}$ Measured radio power at 1.4 GHz (lower limit) Fitted Radio power at 1.4 GHz
mJy mJy $\rm 10^{25}W\,Hz^{-1}$ $\rm 10^{25}W\,Hz^{-1}$
@@
Abell 2142 $26\pm 2$ $44\pm 5$ $05\pm 0.01 0.0$$08\pm 0.01$
Notes: Column 1: name of the cluster hosting the halo; Column 2: Measured flux density of the halo at 1.5 GHz for Abell 2744 and MACS J0717 and 1.28 GHz for PLCK G287.0+32.9, Bullet, and Abell 2142; Column 3: Fitted flux density of the halo, Col. 4: Radio power obtained from flux density reported in Col 2; Column Radio power obtained from flux density reported in Col 3.
Figure 12: Radial profiles of the halos in Abell 2142 (left) and PLCK G287.0+32.9 (right). Profiles fitted with a single-component exponential model are shown with dashed lines. The used annuli are shown in Figure 7 Left: Radial profile of the Abell 2142 halo at 144 MHz, extracted from an image where unrelated sources were masked based on the LOFAR HBA map (red data points) and from an image where unrelated sources were masked using the sensitive 1.3 GHz MeerKAT map (lightcoral data points). The blue data points/line show the residual radial profile obtained by subtracting the uvuv-subtracted sources profile from the masked out sources profile. Top right: PLCK G287.0+32.9 halo profiles at 815 MHz and 350 MHz obtained after subtracting discrete sources from the $uv$-data. Both plots exhibit a second outer component resulting from the incomplete subtraction of discrete sources.
@@ 4.3 Radio emissivity and radio power
Radio halos are observed to exhibit a relatively narrow distribution of radio emissivity, with a characteristic value of approximately $\rm 10^{-42}\,ergs^{-1}\,cm^{-3}\,Hz^{-1}$ (e.g., Murgia et al., 2009). This is based on the assumption of a homogeneous/filling $\rm factor=1$ emitting volume. The emissivity of mega-halos (second component) is reported to be approximately $10^{-44}\,\rm ergs^{-1}\,cm^{-3}\,Hz^{-1}$, about a factor of 20-25 lower than that of classical radio halos (Cuciti et al., 2022).
@@
[…]
We find that the observed halos display a comparable emissivity of approximately $10^{-42}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$. In Figure 11, we present the $I_{0}-r_{e}$ plane at 1.4 GHz for our sample (red points), along with a comparison to a sample of 14 known halos from (Murgia et al., 2009; Vacca et al., 2011; Murgia et al., 2024). We note that Murgia et al. (2024) also masked unrelated sources, making the results directly comparable. Despite their different LLS, we find that the emissivities of halos with LLS>2>2 Mpc are remarkably similar to those of classical halos (Murgia et al., 2024). Moreover, the e-folding radius of these five halos is only about 50 kpc larger than the average e-folding radius reported by Murgia et al. (2024)
The halo in Abell 2142 seems different as its emissivity is about a factor of 10 lower than the other halos with LLS>2>2 Mpc, namely $\langle J\rangle=(1.43\pm 0.09)\times 10^{-43}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$ (estimated using an integrated spectral index of 1.2-1.2, i.e., mean spectral index of the three components). The prototype radio halo in Coma has the lowest emissivity observed so far at 1.4 GHz, with $\langle J\rangle\simeq 5\times 10^{-44}\,\mathrm{erg\,s^{-1}\,cm^{-3}\,Hz^{-1}}$ (Murgia et al., 2024). We note that the Coma and Abell 2142 halos are peculiar, although both are nearby clusters and their radii are comparable to other halos with large extents, their central brightnesses are about an order of magnitude fainter (see Figure 11). The lower emissivities of the Abell 2142 and Coma halos suggests that halos with such low emissivity may not be uncommon.
@@
where $D_{L}$ is the luminosity distance to the cluster, $S_{\nu}$ is the flux density at a frequency $\nu=1.4$ GHz, and $\alpha$ is the spectral index used in the $k$-correction. For the measured flux density of these halos from 1.5 or 1.38 GHz VLA or MeerKAT maps, see Table 7. We emphasize that flux densities were measured within the halo region where the signal exceeds the $\geq 3\sigma_{\rm rms}$. The derived radio powers, at 1.4 GHz, are $P_{\rm PLCK\,G287}=(1.3\pm 0.1)\times 10^{25}\,\mathrm{W\,Hz^{-1}}$ (using $\alpha=-1.15$), $P_{A2744}=(1.4\pm 0.1)\times 10^{25}\,,\mathrm{W\,Hz^{-1}}$ (using $\alpha=-1.15$; Rajpurohit et al., 2021a; Pearce et al., 2017), $P_{\rm MACS\,J0717}=(2.2\pm 0.3)\times 10^{25}\,\mathrm{W\,Hz^{-1}}$ (using $\alpha=-1.5$; Rajpurohit et al., 2021b), $P_{\rm Bullet}=(2.7\pm 0.4)\times 10^{25}\,\mathrm{W\,Hz^{-1}}$ (using $\alpha=-1.1$), $P_{A2142}=(5.2\pm 1.3)\times 10^{23}\,\mathrm{W\,Hz^{-1}}$ (using $\alpha=-1.2$; Riseley et al., 2024). These values represent lower limits, as part of the halo emission is excluded/masked due to the presence of discrete sources.
Figure 14: Left: Sectors used to extract the PLCK G287.0+32.9 halo radial profiles shown in Figure  15 middle panels. Right: Different annuli centers (marked with "+" symbol and a circle of radius 10″) used to extract radial profiles shown in Figure  15, bottom panels. These sectors are overlaid on the 50″MeerKAT UHF uv-subtracted point source image.
@@
$S=2\pi f\cdot r_{e}^{2}I_{0}\quad\text{(mJy)},$ (4)
where $r_{e}$ and $I_{0}$ are in units of arcseconds and $\rm mJy\,arcsec^{-2}$, respectively, and $f\approx 0.8$ accounts for the enclosed flux within $3r_{e}$. The fitted flux densities and the corresponding radio powers are summarized in Table 7. The fitted and measured flux densities of the radio halos in Abell 2744, MACS J0717+3745, and the Bullet Cluster are consistent. However, the fitted flux densities for the halos in PLCK G287.0+32.9 and Abell 2142 are lower than the measured values by a factor of two. Murgia et al. (2009) reported that a tight relation exists between the total radio power of halos and e-folding radius, i.e., $P_{1.4}\propto r_{e}^{3}$. The halos analyzed in this work are consistent with this relation and align well with known classical halos. Moreover, powerful radio halos are expected to be more extended, corresponding to larger e-folding radii (Cassano et al., 2007; Murgia et al., 2009). This also implies that powerful halos should be detectable out to larger radii.
The lower limits and the fitted radio power imply that, excluding the halo in Abell 2142, the remaining systems lie above the established scaling relation between cluster mass and radio power observed for halos at 1.4 GHz. In contrast, the halo in Abell 2142 lies below this relation, as reported by Riseley et al. (2024); Bruno et al. (2023). This also implies that radio powers of halos reported in the literature may be either underestimated (Shimwell et al., 2014; Bonafede et al., 2014) or overestimated (van Weeren et al., 2009; Bonafede et al., 2009; Venturi et al., 2017) due to poor sensitivity and resolution. This in turn impacts the established $P_{\rm 1.4\,GHz}-M_{500}$ scaling relations of radio halos.